What proportion of each source of capital does DragonFlights use?What is the cost of equity for DragonFlights?What is the cost of debt for DragonFlights?DragonFlights, Inc. is planning on purchasing a new flying dragon for their new route to Volantis. The cost of the dragon is
Answer #1

Q1:

Value of equity E =share price*no of shares =116*1 million =116 million

Value of debt D=no of bonds * bond price =50000*1250 = 62.5 million

Total value V= 116+62.5 =178.5

Proportion of equity = E/V =116/178.5 =64.98%

Proportion of Debt =D/V= 62.5/178.5 =35.01%

Q2:

cost of equity = risk free rate+ beta*market risk premium

=2.5+2.8*7 =22.1%

Q3:

cost of debt =YTM=10.9%

cost of debt is the YTM of the bond which can be calculated as below;

Bond cashflows
Years 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26 27 28 29 30 31 32 33 34 35 36 37 38 39 40
Price 1250
Coupon payment 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70
Par value 1000
Total cashflows -1250 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 1070
IRR 5.45%
YTM= 2*IRR 10.90%